OpenStudy (anonymous):

X & Y vary inversely. 1) Inverse variation equation when x=-3 and y=8. 2) Also, what is y when x= 6?

OpenStudy (mathstudent55):

Direct variation is: y = kx Inverse variation is: y = k/x Since you have y = 8 at x = -3, we can find k: 8 = k/-3 k = -24 Our inverse variation equation is: y = -24/x Now find y when x = 6: y = -24/6 y = -4

OpenStudy (anonymous):

Thanks so much !

OpenStudy (mathstudent55):

wlcm